OpenStudy (anonymous):

A ballon has a circumference of 28 in use the circumference to approximate the surface area of the ballon to the nearest square inch

OpenStudy (festinger):

Seeing that no one has given an answer this is my attempt: The circumference is \[28=2\pi r\] Rearranging yields \[r=\frac{14}{\pi}\] The surface area for a sphere is: \[S=4\pi r^{2}\] Substituting \(r=\frac{14}{\pi}\) gives \[S=4\pi \frac{14}{\pi} \frac{14}{\pi}=\frac{784}{\pi}\]